OpenStudy (anonymous):

indefinate integral of 1/xln(3x) dx

OpenStudy (anonymous):

By using substitution method put t = In(3x) then diff both sides and substitute it in the eq

OpenStudy (anonymous):

That is where I get stuck. When differentiating 1/xt now,,, isn't that now \[1/x \times 1/t\] when differentiates to \[lnx \times lnt\]???